OpenStudy (anonymous):

3/4b- 1/6= 1/2 solve for b

jigglypuff314 (jigglypuff314):

add 1/6 to both sides multiply both sides by 4/3

jigglypuff314 (jigglypuff314):

3/4b - 1/6 = 1/2 + 1/6 +1/6 3/4 b = 4/6 3/4 b = 2/3 \[b = \frac{ 2 }{ 3 } * \frac{ 4 }{ 3 }\] multiply and that's your answer :)

OpenStudy (anonymous):

thanks